OpenStudy (anonymous):

Challenge! Ready? A tree that is 10 yards tall casts a shadow 14 yards long. Find the angle of elevation from the tip of the shadow to the top of the tree. Round to the nearest degree.

OpenStudy (wolf1728):

The shadow length would be the adjacent side and the tree would be the opposite side. tangent (angle) = opp/adj tangent (angle) = 10/14 tangent (angle) = 0.71429 Can you figure it out from here?

OpenStudy (anonymous):

I already know the answer. It is 36 degrees

OpenStudy (wolf1728):

okay the arc tangent of .71429 is 35.538 Degrees Which rounds to 36 degrees Did you want to know how it is done?

OpenStudy (anonymous):

I know how it is done, thank you. The answer is what I need confirmation on.

OpenStudy (wolf1728):

Okay - well there it is :-)

OpenStudy (anonymous):

Thanks, I just wasn't sure.

OpenStudy (wolf1728):

Okay and thanks for the medal :-)